OpenStudy (anonymous):

Find f^-1(x) f(x)=3sqrtx-7 Please help

hartnn (hartnn):

let y= 3\(\sqrt x\)-7 interchange, x and y x= 3\(\sqrt y\)-7 now can u fin y in terms of x ?

OpenStudy (anonymous):

no I have no clue what I'm doing and really need help

hartnn (hartnn):

to find inverse function, u find x in terms of f(x) for convenience, i took it as y when i interchanged variables, now i need to find y in terms of x, so can u ?

OpenStudy (anonymous):

I can't do math for the life of me and can't figure it out

OpenStudy (anonymous):

Is it (x+7)^2 this only thing that I can come up with

hartnn (hartnn):

oh, u get (x+7)^2 = 9y , isn't it ?

OpenStudy (anonymous):

I have no idea what the correct answer is, I don't understand where your coming up with the 9y

hartnn (hartnn):

let me show you x= 3 \(\sqrt y-7\) Adding 7 to both sides x+7 =3 \(\sqrt y\) squaring both sides (x+7)^2=9y got this ?

OpenStudy (anonymous):

so f^-1(x)=(x+7)^2= 9y ?? Understanding that x+7 ^2 = (x+7) and 3^2 = 9

hartnn (hartnn):

to get f^-1 x, u need to isolate y so y= (x+7)^2/9 .....(dividing 9 on both sides) so that will be your f^-1(x) = (x+7)^2/9 got it ?

OpenStudy (anonymous):

Yes I do thanks you

hartnn (hartnn):

welcome :)